A long-distance runner runs 2 miles south and then 8 miles east. How far is the runner from the starting point? Round your answer to the nearest thousandth if necessary
Answer by Boreal(15235)

these are two legs of a right triangle.
the hypotenuse is the distance from the starting point
sqrt(hypot)=sqrt (2^2+8^2)=sqrt(68) or 2 sqrt (17) miles
This is 8.246 miles